OpenStudy (anonymous):

if f(8)=5 and f'(8)=12, find the best estimate for f(9).

OpenStudy (anonymous):

8/5=9/x I guess

OpenStudy (anonymous):

you guess?

OpenStudy (mathstudent55):

How can a function have f(8)= 5 and f(8) = 12. This is not a function then.

OpenStudy (anonymous):

@mathstudent55 it says f prime(8)= 12

OpenStudy (mathstudent55):

Oh, ok, thanks. I couldn't see that.

OpenStudy (anonymous):

@Hero, what do you mean, how would I use it?

hero (hero):

I was mainly speaking against @timo86m setup

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so, what's next?

OpenStudy (anonymous):

ok, make me think :)

OpenStudy (anonymous):

isn't it (8,5)?

OpenStudy (anonymous):

ok, so you don't have to use f '(8)=12 at all?

OpenStudy (anonymous):

oh ok, thank you!

OpenStudy (anonymous):

would it be the same if i were to use the point slope formula, and then plug 9 into the equation I get out of it?

hero (hero):

Okay, I have it for sure. Sometimes I have to talk my way through these. Basically the slope is given: f'(x) = m In this case: f'(8) = 12 Therefore the slope is 12. Thus rise/run = y/x = 12/1 Now, plot the given point (8,5). Next, move the point up 12 units, and 1 unit to the right. The next point is (9, 17)

OpenStudy (anonymous):

ok thank you :)

OpenStudy (mathstudent55):

@hero Does a problem like this assume that f(x) is a linear function?

hero (hero):

We're not assuming the function is linear. We used the definition of slope to find the next point.

OpenStudy (mathstudent55):

But the slope provided was at the specific point where x = 8. Couldn't the slope also have changed between x = 8 and x = 9? If f(x) is a third degree polynomial, the slope would be quadratic, and may be different at f(9).

hero (hero):

By the way, it says to estimate. So we can assume the that the next point is close to (8,5)

OpenStudy (mathstudent55):

Ok, I think that's the key.

OpenStudy (mathstudent55):

Thanks

hero (hero):

There's no other way to figure this out

hero (hero):

What you mentioned about degree of polynomial is exactly the reason why there is no other way. We are not given the formulas for f(x) or f'(x). So we have to make some assumptions in order to solve it.

OpenStudy (mathstudent55):

I get it now, thanks.